# BSTcontrol

### v2

@@ 4.3.2 Spatial AirFusion for Integrated Sensing and Edge AI
The integration of distributed sensing capabilities with edge AI functionalities (ISEA) often involves deploying complex, potentially large perception models at the network edge for applications like autonomous driving or collaborative robotics [190]. These applications rely on fusing spatial data (e.g., high dimensional voxel features from LiDAR or camera sensors) from multiple agents to build a comprehensive environmental understanding. Transmitting these voluminous spatial features from numerous agents to a fusion center presents significant communication bottlenecks, hindering real time performance crucial for safety critical systems. Spatial AirFusion is a specialized AirComp framework designed specifically to overcome this challenge by exploiting the unique characteristics of spatial sensing data generated for and processed by these AI models. Its key innovation lies in leveraging spatial feature sparsity (a property often observed in the intermediate representations within large perception models, where many spatial locations might be inactive) and broadband channel frequency diversity. The protocol involves agents reporting sparsity patterns, followed by the fusion center performing optimized resource allocation. This process intelligently maps voxels to specific subcarriers and allocates power to maximize the minimum receive signal to noise ratio, ensuring reliable aggregation of the features needed by the central inference part of the large AI model. By tailoring AirComp to the spatial domain and considering the specific data structures relevant to AI perception models, Spatial AirFusion significantly improves sensing performance and reduces aggregation errors compared to generic AirComp, directly benefiting the accuracy and efficiency of distributed large AI models in ISEA contexts.
@@ 4.3.3 AirPooling for Generalized Function Computation in Large AI Models
Standard AirComp is inherently suited for computing nomographic functions, typically involving summations [11, 225]. However, many large AI models, particularly deep neural networks, employ a wider variety of aggregation operations within their architectures. Operations like Max-Pooling, common in CNNs for feature downsampling, or other non linear aggregation functions used in multi view sensing or attention mechanisms, are not directly computable via basic AirComp. AirPooling extends AirComp capabilities to efficiently realize such crucial non nomographic functions over the air, directly supporting the distributed implementation or inference of large AI models containing these operations [191]. It specifically addresses Average-Pooling and Max-Pooling by utilizing the mathematical properties of the generalized p-norm function, which smoothly interpolates between summation and the maximum function. AirPooling implements this by decomposing the target computation into appropriate preprocessing at the edge devices and postprocessing at the server, leveraging the underlying summation provided by the AirComp MAC. The design optimizes a configuration parameter to carefully balance the accuracy of approximating the desired AI specific function against the noise amplification inherent in AirComp. Task oriented optimization, connecting the AirPooling error to final inference accuracy (e.g., using classification margin theory), ensures that the computation directly benefits the performance of the large AI model. AirPooling, therefore, significantly broadens the applicability of AirComp, enabling the efficient wireless execution of essential computational building blocks found within diverse large AI architectures in distributed settings.
4.4 Physical Layer Security for Large AI Model
PLS is gaining recognition as a vital component for securing large AI models within 6G communications. Traditional cryptographic methods face challenges in meeting the stringent latency and scalability demands of emerging 6G applications like mMTC and URLLC, as highlighted by [128]. PLS offers a complementary approach by leveraging the inherent randomness of the wireless channel to establish secure communication directly at the physical layer, thus providing a lightweight yet robust security mechanism. This is particularly crucial as 6G networks are envisioned to support increasingly intelligent and automated services relying on large AI models, necessitating enhanced security measures from the ground up.
To effectively secure large AI models, PLS techniques such as secret key generation and wiretap channel coding are essential. It is demonstrated that the potential of integrating AI with PLS to significantly improve security performance [192]. AI enhanced PLS schemes can achieve superior SNR performance and ensure near perfect secrecy for legitimate users even in complex fading environments. Furthermore, advanced techniques like space time coding based PLS offer substantial security gains and improved reliability, making them suitable for securing the data intensive and mission critical applications of large AI models in 6G.
Looking ahead, PLS is poised to play an increasingly important role in the broader 6G security landscape. As networks become more context aware and require adaptive security levels, the inherent flexibility of PLS becomes a key advantage. It is emphasized that PLS can provide information theoretic security guarantees with lightweight mechanisms and can be integrated with traditional cryptography in hybrid protocols for enhanced protection [128]. This hybrid approach, combining the strengths of both cryptographic and physical layer techniques, is likely to be a crucial direction for achieving comprehensive and adaptable security solutions for 6G and the secure deployment of large AI models.
@@ 5 Emerging Technologies for WLAM
Hyper-dimensional Computing [229] Execute language classification tasks efficiently and robustly. Introduce a hardware architecture for a classifier based on hypervector. Enhanced expressive power
[230] Categorizing news articles based on a continuous stream of input letters for text classification. Demonstrate a software classification framework employing hyperdimensional computing.
[231] Execute cognitive tasks efficiently with constrained energy budget and computing resources. Present a overview on the paradigm, algorithms, and applications of hyperdimensional computing.
Quantum Computing [232] Develop wireless communication systems that are fast, reliable, secure and energy-efficient. Discuss the quantum algorithms to improve the physical and network layers of wireless communications. Powerful computing capability and improved security
[233] Latency-sensitive and computing-intensive tasks with many quantum computers. Introduce the concept of distributed quantum computing and its applications.
[234] Integrate quantum computing with machine learning for wireless communications. Discuss the state-of-the-art quantum machine learning algorithms and potential applications in wireless communications.
[235] Resource allocation in wireless communication environments. Present a novel quantum neural network and a reinforcement-enhanced version.
[236] Distributed resource optimization in wireless communication systems. Present a federated quantum neural network framework utilizing quantum teleportation.
[237] Natural language processing with quantum neural networks. Propose a novel deep neural network model with entanglement embedding module.
Physical Inspired Neural Networks [238] Fast mapping between codes and radiation patterns for electrically large meta-surfaces in beamforming. Propose PINN for code-to-pattern mapping and PINN-guided DNN for pattern-to-code mapping, combined for intelligent beamforming. More responsive, adaptable, and explainable system
@@
[241] MU-MIMO beamforming scenarios in dynamic wireless environments with noise interference. Introduce a gradient-based liquid neural network framework to effectively perform beamforming.
[242] Beam tracking leveraging mmWave to predict best beam index for moving users. Introduce a robust beam tracking framework employing multi layers of liquid neurons.
Hyper-Networks [243] Allocate resource in RIS-assisted communication systems with deep neural networks. Propose a hypernetwork-based approach to generate the beamforming vectors conditioned conditioned on the input user weights. Quick adaptation and tailoring
[244] Develop AI models specifically customized to match the user’s provided data or task descriptions. Propose a hypernetwork-based framework to rapidly generate customized AI models.
Next-Gen Sequence Modeling Networks [245] Long sequence modeling across various modalities (text, audio, picture) where Transformers are inefficient. Propose Mamba architecture based on selective SSMs with hardware-aware parallel algorithm and simplified design. Improved inference speed and longer context
[246] General sequence processing tasks aiming for efficiency and performance with linear scaling. Propose RWKV combining Transformer-style parallel training and RNN-style efficient inference using linear attention.
[247] Long context sequence modeling needing linear complexity and expressive hidden states. Propose custom layers with self-supervised learning updated hidden states.
Emerging technologies are pivotal for the evolution of WLAM, enabling transformative capabilities in wireless communications. This section delves into these technologies, categorizing them into emerging computing paradigms and emerging neural network architectures. Table 5 summarizes related research. Through this exploration, we highlight the profound impact of these technologies in shaping the future of WLAM.
@@ 5.1.1 Hyperdimensional Computing
HDC is an emerging technology inspired by the remarkable representation abilities of high-dimensional vector spaces. Unlike traditional computing methods that use low-dimensional data representations, HDC leverages high-dimensional vectors, often comprising thousands of dimensions, to encode information efficiently [248]. The basic HDC pipeline includes three stages: encoding, training, and comparison [249]. At the encoding stage, inputs are mapped into high-dimensional vectors, which are robust to noise and capable of representing complex data structures. During training, these vectors are stored in the associative memory, grouping similar inputs into classes. In the comparison stage, query vectors are matched against stored classes by measuring similarity, typically using the Hamming distance. In the following part we delve into the advantages, various applications, and future prospects of HDC.
@@
HDC has been successfully applied in areas like language recognition [229], text categorization [230], speech recognition [250], etc. It offers several advantages with respect to the implementation in AI [231], including significantly lower power consumption and latency compared to traditional deep neural networks, making it suitable for edge computing devices with limited resources. Besides, the high-dimensional nature of HDC provides robustness against noise and uncertainty that is beneficial for real-world applications where data may be incomplete or noisy. Furthermore, the ability of HDC to handle a vast number of unique vectors enhances its scalability across various applications.
@@
Although promising for the benefits mentioned above, HDC is still in its early stages when it comes to its application in large AI models [251]. The well-known transformer architecture, which is central to many popular AI models, could significantly benefit from the integration of HDC by utilizing much larger embedding dimensions. High-dimensional vectors in HDC offer richer and more expressive representations, which can enhance the capabilities and performance of AI models, potentially benefiting applications in RL environments.

### v4

@@ 4.3.2 AirComp for Large Multi-Modal Models in ISAC
The rise of large multi-modal AI models for applications like autonomous driving and collaborative robotics has created new challenges for ISAC. These large AI models often rely on fusing high-dimensional spatial data (e.g., voxel features from LiDAR or camera sensors) from multiple distributed agents to build a comprehensive environmental understanding [liu2024over]. Transmitting these voluminous features to a central fusion node via conventional digital methods presents a significant communication bottleneck, hindering the real-time performance critical for such safety-sensitive systems. Spatial AirFusion is a specialized AirComp framework designed to overcome this challenge. It exploits the unique characteristics of spatial sensing data, such as feature sparsity often found in the intermediate representations of large perception models. By intelligently mapping features to subcarriers and optimizing power, it enables reliable over-the-air aggregation, directly benefiting the accuracy and real-time viability of distributed large multi-modal models in ISAC contexts.
@@ 4.3.3 AirPooling for Generalized Function Computation in Large AI Models
Standard AirComp is inherently suited for computing nomographic functions, typically involving summations [wang2024over, zhu2021over]. However, many large AI models, particularly deep neural networks, employ a wider variety of aggregation operations within their architectures. Operations like Max-Pooling, common in CNNs for feature downsampling, or other non-linear aggregation functions used in multi-view sensing or attention mechanisms, are not directly computable via basic AirComp.
To address this limitation, AirPooling extends AirComp capabilities to efficiently realize such crucial non-nomographic functions over the air, directly supporting the distributed implementation or inference of large AI models containing these operations. It specifically addresses Average-Pooling and Max-Pooling by utilizing the mathematical properties of the generalized p-norm function, which smoothly interpolates between summation and the maximum function. AirPooling implements this by decomposing the target computation into appropriate pre-processing at the edge devices and post-processing at the server, leveraging the underlying summation provided by the AirComp mechanism [liu2023over].
@@
The design optimizes a configuration parameter to carefully balance the accuracy of approximating the desired AI-specific function against the noise amplification inherent in AirComp. Task-oriented optimization, connecting the AirPooling error to final inference accuracy (e.g., using classification margin theory), ensures that the computation directly benefits the performance of the large AI model. AirPooling, therefore, significantly broadens the applicability of AirComp, enabling the efficient wireless execution of essential computational building blocks found within diverse large AI architectures in distributed settings.
4.4 Physical Layer Security for Large AI Models
The deep integration of large AI models into 6G communications fundamentally elevates the importance of security. It creates new attack surfaces and high-value targets. Traditional cryptographic methods, while essential, may face challenges in meeting the stringent latency and scalability demands of distributed AI systems. PLS emerges as a crucial complementary paradigm. It leverages the inherent properties of the wireless channel to provide a lightweight and robust layer of defense for the operations of large AI models [mitev2023physical].
4.4.1 Fundamental Limits of Secure Federated Learning
Securing the federated learning process for large AI models over wireless channels introduces a complex interplay of requirements. The system must satisfy the privacy definitions of federated learning. It must also meet the strict confidentiality demands of physical layer security, where information leakage to an external eavesdropper should be negligible. This creates a fundamental trade-off. The work in [infocomfederated] addresses this challenge by proposing a novel architecture for wireless federated learning that incorporates both a privacy mechanism and secure channel coding. The authors formally characterize the theoretical performance boundaries by deriving the capacity-equivocation region. This analysis establishes the fundamental limits on securing the high-dimensional parameter updates that are characteristic of large AI models.
4.4.2 Resilience Against Physical Layer Attacks
Beyond theoretical limits, the practical performance of a large AI model is directly impacted by the quality of the physical layer. Adversarial attacks like jamming can corrupt the transmission of critical information. This includes the embeddings or gradients exchanged during split federated learning. The work in [federatedPHY] provides a rigorous analysis of this problem. It establishes a direct mathematical connection between the training loss error of a large language model and the communication mean squared error (MSE) over the wireless link. This insight formally demonstrates why the physical layer matters for the convergence of distributed learning. Building on this analysis, the authors propose a resilience-by-design framework. This framework includes methods for beamforming and resource allocation that ensure high task performance and model convergence even under worst-case jamming conditions.
4.4.3 Lightweight Security for Scalable Architectures
The distributed nature of many large AI model applications involves a massive number of participating devices. Many of these devices may be resource-constrained. Deploying and managing complex cryptographic protocols across thousands of heterogeneous nodes presents a significant overhead. Physical layer security offers a lightweight solution well-suited for such scalable systems [mitev2023physical, 10850506]. It exploits existing physical channel properties to provide security without requiring extensive computational resources. This low-latency and low-overhead nature makes PLS an ideal component of a defense-in-depth strategy. It can work in concert with traditional cryptography to build a holistic and efficient security architecture for trusted and autonomous networks driven by large AI models in 6G and beyond.
4.4.4 Illustrative Privacy and Security Threats and PLS Contributions
To illustrate the threats that necessitate the PLS solutions described above, we analyze specific scenarios for both privacy and security, detailing the direct contributions of PLS in mitigating them.
Eavesdropping Attack
A primary threat to privacy in distributed learning is posed by inference attacks through passive eavesdropping. Consider a wireless hierarchical federated learning system where an edge server transmits aggregated gradient updates to a cloud server [Zhang2023Finite]. The attack process involves an eavesdropper capturing the signal $Z_{i}(t)$ at each time instant $i$ of communication round $t$. The received signal at the eavesdropper can be modeled as:
$Z_{i}(t)=gX_{i}(t)+\tilde{g}\tilde{X}_{i}(t)+\eta_{e,i}(t),$ (27)
where $X_{i}(t)$ and $\tilde{X}_{i}(t)$ are the legitimate uplink and downlink signals, $g$ and $\tilde{g}$ are the complex channel gains to the eavesdropper, and $\eta_{e,i}(t)$ is additive noise. The consequence is a direct breach of privacy, as the eavesdropper can analyze the captured signals to infer properties of the training data. To counter this, PLS contributes by designing secure channel coding schemes. The achievable secure rate $R_{t}$ for a given communication round $t$ can be quantified, for example, by the bound:
$R_{t}=\frac{1}{N_{t}}\log\left(\frac{3\text{SNR}|h|^{2}}{[Q^{-1}(\frac{\tau}{8})]^{2}}\left(1+\frac{\text{SNR}|h|^{2}}{\Psi_{1}\Psi_{2}}\right)^{N_{t}-1}\right).$ (28)
Here, $R_{t}$ is the rate in bits per channel use, $N_{t}$ is the blocklength, SNR is the signal-to-noise ratio of the legitimate link with channel gain $h$, $Q^{-1}(\cdot)$ is the inverse Gaussian Q-function related to the decoding error probability $\tau$, and $\Psi_{1},\Psi_{2}$ are terms related to the feedback channel quality. This formula demonstrates that PLS can provide a quantifiable positive secure rate, ensuring confidentiality against the eavesdropper.
Jamming Attack
On the security front, a direct threat to system integrity is jamming for model performance degradation. In an SFL system fine-tuning an LLM, an adversary can execute a worst-case jamming attack during the uplink [federatedPHY]. The received signal at the base station is modeled as a superposition of the legitimate signals $x_{qnk}$ and a jamming signal $z_{nk}$:
$y_{nk}=\sum_{q\in\mathcal{Q}}H_{qnk}x_{qnk}+z_{nk},$ (29)
where $z_{nk}=G_{nk}u_{nk}+\eta_{nk}$ consists of the jamming signal $u_{nk}$ through the jammer channel $G_{nk}$ and receiver noise $\eta_{nk}$. The consequence is catastrophic, as the jamming directly increases the mean squared error of the communication, causing the server to train the model on corrupted embeddings and preventing convergence. PLS directly addresses this by enabling resilient-by-design communication. The framework in [federatedPHY] proposes designing receive filters based on a surrogate interference covariance matrix that robustly accounts for the jammer:
$\tilde{C}_{z_{nk}}=\sigma^{2}I_{N_{R}}+\eta A_{B}(\theta_{G})A_{B}^{H}(\theta_{G}),$ (30)
where $\tilde{C}_{z_{nk}}$ is the constructed surrogate matrix, $\sigma^{2}$ is the noise power, $I_{N_{R}}$ is the identity matrix for $N_{R}$ receive antennas, $\eta$ is a resilience hyperparameter, and $A_{B}(\theta_{G})$ is the array manifold based on the known directions of arrival $\theta_{G}$ of the jammer. This allows the system to safeguard the integrity of the training process.
Man-in-the-Middle Attack
Another critical security threat targets authentication and confidentiality via man-in-the-middle attacks, particularly in protocols that generate secret keys from channel randomness [mitev2023physical]. The attack mechanism involves an adversary transmitting a pre-coded signal $W$ to two legitimate devices, Alice and Bob, during their channel estimation phase. The resulting observations are modeled as:
$Z_{A}=XH+W+N_{A},$ (31)
$Z_{B}=XH+W+N_{B},$ (32)
where $Z_{A}$ and $Z_{B}$ are the signals received by Alice and Bob, $X$ is the legitimate probe signal over the reciprocal channel $H$, $N_{A}$ and $N_{B}$ are receiver noise, and $W$ is the malicious signal. The consequence is a complete breakdown of confidentiality, as the shared randomness is contaminated by the attacker’s known signal $W$. The contribution of PLS lies in redesigning the physical layer protocol. A countermeasure described in [mitev2023physical] involves both users transmitting independent, randomized probe signals, $X$ from Alice and $Y$ from Bob. After post-processing their received signals, their new observations become:
$\tilde{Z}_{A}=XYH+XW+XN_{A},$ (33)
$\tilde{Z}_{B}=XYH+YW+YN_{B},$ (34)
where $\tilde{Z}_{A}$ and $\tilde{Z}_{B}$ are the processed observations. In this new formulation, the shared randomness term is $XYH$, while the attacker contributions $XW$ and $YW$ are no longer identical. This physical layer procedure makes the signal of attacker uncorrelated with the shared secret component, preserving the integrity of the key generation process.
4.5 Practical Framework for Enabling Large AI Models
The successful deployment and operation of large AI models within distributed wireless environments require a robust framework that systematically addresses computation, communication, and security. This framework provides a guideline for practitioners to build efficient and reliable systems.
4.5.1 Deployment Strategy Selection
The first step is deployment strategy selection, which involves choosing the appropriate distributed collaborative machine learning paradigm [kairouz2021advances]. Depending on data distribution, privacy requirements, and device capabilities, a practitioner might select FL for parallel training on decentralized data [10436365], SL for scenarios with resource constrained clients [10843784], or FSL as a hybrid approach that combines the benefits of both [thapa2022splitfed]. This foundational choice determines the overall system architecture and the nature of interactions between nodes.
4.5.2 Resource and Model Partitioning
The second phase is resource and model partitioning. Once a distributed strategy is chosen, the next critical task is to allocate resources and partition the model across the network, which is especially important for edge devices with limited computing power [lin2023pushing]. For heterogeneous systems, this involves the intelligent assignment of model components. For example, in a MoE model, the expert networks can be distributed across multiple mobile devices while a central server manages the gating network [wdmoe]. Alternatively, for other architectures, tensor parallelism can be used to partition the internal weight matrices of a large model across different devices to enable collaborative inference [zhang2025communication]. Successfully solving these joint allocation and partitioning problems is central to realizing the full potential of collaborative edge intelligence.
4.5.3 Communication Efficiency Optimization
In the third phase, the framework must address communication efficiency optimization. The exchange of high dimensional gradients or activations is a major bottleneck in distributed AI. This framework should incorporate advanced wireless techniques to mitigate this challenge. A key technology is AirComp, which leverages the superposition property of wireless channels to perform fast data aggregation [wang2024over, zhu2021over]. In the context of federated learning, this approach, known as Air FEEL, enables one shot aggregation of model updates, significantly enhancing communication efficiency and making the distributed training of large models more feasible [cao2024overview].
4.5.4 Security Integration
The final and essential component of the framework is security integration. The distributed nature of these systems introduces new vulnerabilities that must be addressed from the design stage. PLS emerges as a crucial complementary paradigm, offering a lightweight yet robust layer of defense for the operations of large AI models [mitev2023physical]. It can provide confidentiality for model updates against eavesdroppers [infocomfederated] and enhance the resilience of the training process against physical layer attacks like jamming [federatedPHY]. Integrating these security measures is paramount to ensuring the integrity and privacy of distributed AI systems, fostering trust in the autonomous networks of 6G and beyond.
@@ 5 Emerging Technologies for WLAM
Hyper-dimensional Computing [rahimi2016robust] Execute language classification tasks efficiently and robustly. Introduce a hardware architecture for a classifier based on hypervector. Enhanced expressive power
[najafabadi2016hyperdimensional] Categorizing news articles based on a continuous stream of input letters for text classification. Demonstrate a software classification framework employing hyperdimensional computing.
[hassan2021hyper] Execute cognitive tasks efficiently with constrained energy budget and computing resources. Present a overview on the paradigm, algorithms, and applications of hyperdimensional computing.
Quantum Computing [botsinis2018quantum] Develop wireless communication systems that are fast, reliable, secure and energy-efficient. Discuss the quantum algorithms to improve the physical and network layers of wireless communications. Powerful computing capability and improved security
[buhrman2003distributed] Latency-sensitive and computing-intensive tasks with many quantum computers. Introduce the concept of distributed quantum computing and its applications.
[narottama2023quantum] Integrate quantum computing with machine learning for wireless communications. Discuss the state-of-the-art quantum machine learning algorithms and potential applications in wireless communications.
[narottama2021quantum] Resource allocation in wireless communication environments. Present a novel quantum neural network and a reinforcement-enhanced version.
[narottama2023federated] Distributed resource optimization in wireless communication systems. Present a federated quantum neural network framework utilizing quantum teleportation.
[chen2021quantum] Natural language processing with quantum neural networks. Propose a novel deep neural network model with entanglement embedding module.
Physical Inspired Neural Networks [li2022intelligent] Fast mapping between codes and radiation patterns for electrically large meta-surfaces in beamforming. Propose PINN for code-to-pattern mapping and PINN-guided DNN for pattern-to-code mapping, combined for intelligent beamforming. More responsive, adaptable, and explainable system
@@
[ncps_beamforming] MU-MIMO beamforming scenarios in dynamic wireless environments with noise interference. Introduce a gradient-based liquid neural network framework to effectively perform beamforming.
[zhu2024robustcontinuoustimebeamtracking] Beam tracking leveraging mmWave to predict best beam index for moving users. Introduce a robust beam tracking framework employing multi layers of liquid neurons.
Hyper-Networks [abouamer2023flexible] Allocate resource in RIS-assisted communication systems with deep neural networks. Propose a hypernetwork-based approach to generate the beamforming vectors conditioned conditioned on the input user weights. Quick adaptation and tailoring
[tang2024modelgpt] Develop AI models specifically customized to match the user’s provided data or task descriptions. Propose a hypernetwork-based framework to rapidly generate customized AI models.
Next-Gen Sequence Modeling Networks [gu2024mamba] Long sequence modeling across various modalities (text, audio, picture) where Transformers are inefficient. Propose Mamba architecture based on selective SSMs with hardware-aware parallel algorithm and simplified design. Improved inference speed and longer context
[peng-etal-2023-rwkv] General sequence processing tasks aiming for efficiency and performance with linear scaling. Propose RWKV combining Transformer-style parallel training and RNN-style efficient inference using linear attention.
[sun2024learninglearntesttime] Long context sequence modeling needing linear complexity and expressive hidden states. Propose custom layers with self-supervised learning updated hidden states.
The communication frameworks discussed in Section 4 establish the necessary foundation for deploying large AI models. However, realizing the stringent performance and efficiency demands of the 6G era requires moving beyond the optimization of existing architectures. This necessitates a paradigm shift, calling for the integration of novel and deeply integrated methodologies to unlock the full potential of WLAMs. Consequently, emerging technologies become pivotal for the evolution of WLAM, enabling transformative capabilities in wireless communications. This section delves into these technologies, organizing them into two distinct categories to provide a clear overview. First, we explore emerging computing paradigms, and then we investigate emerging neural network architectures. Table 8 summarizes related research. Through this exploration, we highlight the profound impact of these technologies in shaping the future of WLAMs.
@@ 5.1.1 Hyperdimensional Computing
HDC is an emerging technology inspired by the remarkable representation abilities of high-dimensional vector spaces. Unlike traditional computing methods that use low-dimensional data representations, HDC leverages high-dimensional vectors, often comprising thousands of dimensions, to encode information efficiently [kleyko2022survey]. The basic HDC pipeline includes three stages: encoding, training, and comparison [thomas2021theoretical]. At the encoding stage, inputs are mapped into high-dimensional vectors, which are robust to noise and capable of representing complex data structures. During training, these vectors are stored in the associative memory, grouping similar inputs into classes. In the comparison stage, query vectors are matched against stored classes by measuring similarity, typically using the Hamming distance. In the following part we delve into the advantages, various applications, and future prospects of HDC.
@@
HDC has been successfully applied in areas like language recognition [rahimi2016robust], text categorization [najafabadi2016hyperdimensional], speech recognition [imani2017voicehd], etc. It offers several advantages with respect to the implementation in AI [hassan2021hyper], including significantly lower power consumption and latency compared to traditional deep neural networks, making it suitable for edge computing devices with limited resources. Besides, the high-dimensional nature of HDC provides robustness against noise and uncertainty that is beneficial for real-world applications where data may be incomplete or noisy. Furthermore, the ability of HDC to handle a vast number of unique vectors enhances its scalability across various applications.
@@
Although promising for the benefits mentioned above, HDC is still in its early stages when it comes to its application in large AI models [ge2020classification]. The well-known transformer architecture, which is central to many popular AI models, could significantly benefit from the integration of HDC by utilizing much larger embedding dimensions. High-dimensional vectors in HDC offer richer and more expressive representations, which can enhance the capabilities and performance of AI models, potentially benefiting applications in RL environments.